OpenStudy (anonymous):

what is the equation of the parabola in vertex form 0=y^2-x+2y-3?

OpenStudy (anonymous):

A=(x+4)=(y+1)^2 B=(x+3)=(y+1)^2 C=(x-3)=(y+2)^2 D=(x+7)^2=(y+2)

OpenStudy (anonymous):

Is the answer C?

OpenStudy (tkhunny):

Why don't you put it in vertex form and see?

OpenStudy (anonymous):

Because i'm having trouble doing so.

OpenStudy (tkhunny):

Okay, what have you tried? 0=y^2-x+2y-3 Move x x = y^2+2y-3 Complete the Square on y x = y^2+2y + ____ -3 - ____ Fill in the blanks. You're almost done.